# Macroscopic States in Bose–Einstein Condensate Dark Matter Model with Axionlike Interaction

### v1

@@ III.2 Analyzing the macroscopic state of BEC DM
Having extracted the model parameters from the galaxy’s rotation‐curve fit (e.g., NGC 2366), we then apply statistical analysis to probe the properties of BEC DM. Specifically, we focus on identifying the macroscopic state of DM in the galaxy. It depends on the fractions of the dense and rarefied phases of DM , and their change may be a phase transition. By mapping all configurations onto a diagram, the state of the galaxy would be represented by a single point there.
@@
When generating a set of solutions $\chi(\xi)$ for different values of $A$ and $\nu$ to support statistical analysis, we ensure that their shapes are similar to the reference solution shown in Fig. 4. To achieve this correspondence, varying $\nu$ at fixed $A$, we choose initial conditions from an “island” similar to that in Fig. 5 and impose certain conditions: $\chi^{\prime}(\xi_{B})>0$ and $\chi(\xi_{B})=\pi$, which should hold simultaneously and define $\xi_{B}$.
Each wave function $\chi(\xi)$ in this set defines a chemical potential $\mu$, an internal pressure $p$, and an internal energy density $e$ according to
@@
Denoting the boundary value of the wave function as $\chi_{B}\equiv\chi(\xi_{B})$ and its derivative as $\chi^{\prime}_{B}\equiv\chi^{\prime}(\xi_{B})$, we proceed to analyze Eq. (78). Integrating by parts the trigonometric expressions, we arrive at
@@
Denoting $\chi_{0}\equiv\chi(0)$ and expanding the set of previously introduced characteristics, we included the mean energy density ${\cal E}$ in Fig. 10(f):
@@
Thus, an increase in quantum fluctuations $\Pi_{\rm q}$ predictably leads to an expansion of the system, see Fig. 10(g). Consequently, the mean densities of particle number and internal energy naturally decrease, as shown in Fig. 10(d) and (f). Additionally, the enhanced fluctuations inhibit the increase of the central particle density, which is proportional to $\chi^{2}_{0}$, see Fig. 10(c). Moreover, the transition to the predominance of the rarefied phase is accompanied by an increase in chemical potential $\nu$, see Fig. 10(a). However, the decreasing part of the chemical potential $u$ in Fig. 10(b) indicates the unstable state. The near-linear dependence on larger $\Pi_{\rm q}$ confirms this.
@@
Valuable information can be also extracted from the mass-radius relation $M(R)$ by associating the total mass $M=4\pi r_{0}^{3}\rho_{0}N$ and the radius $R=r_{0}\xi_{B}$ with dimensionless quantities in Fig. 11(a). Although the number $N$ in the case of the unstable rarefied phase dominance tends to increase rapidly with radius $\xi_{B}>8$, we note the presence of a minimum $N_{\min}$ for relatively dense and compact objects at $\xi_{B}<8$. The existence of a minimum point separating the stable and unstable branches in our approach reproduces the predictions for the axiton BZ19 ; KT94 ; ax5 .
@@
The BEC DM state of NGC 2366, marked by a red dot in Fig. 10 and 11 at $\Pi_{\rm q}\simeq 5.852$, is characterized by the presence of $x\simeq 19.2\%$ of the dense DM in Fig. 10(h), which is concentrated in $4.7\%$ of the volume. According to our model, it belongs to the (blue) continuous trajectory, which is classified as supercritical fluid with respect to the variations in $(A,\Pi_{\rm q})$. However, this conclusion relies on spatially averaged values and overlooks the inhomogeneous DM distribution, as evidenced in Fig. 4 and 9. In addition, the phase trajectory of a macroscopic system can be significantly altered by random or external influences. In fact, BEC DM of NGC 2366 exhibits a structured, fruitlike composition, consisting of a dense core, sparse pulp, and crust. These zones are defined by distinct radial intervals: 1) $r\leq 2.83$ kpc, including a central region for $r\leq 0.94$ kpc with a density ratio $\chi_{0}^{2}/\sigma\simeq 6.04$; 2) $2.83\leq r\leq 5.55$ kpc; 3) $5.55\leq r\leq 7.79$ kpc, representing the outer layer in the model.
@@
On the other hand, a sharp decrease in $\Omega$ signals a fall into an unstable state.
@@
In Fig. 13(b), it is shown that the impact of surface gravity, expressed as ${\rm g}=AN/\xi^{2}_{B}$, also confirms the presence of a first-order phase transition in the model, although ${\rm g}$ is not a thermodynamic function (potential) used to extract critical parameters. This confirms that its value becomes larger as $x$ increases.
@@ IV Conclusion
We developed a statistical approach to the stationary nonrelativistic BEC DM model with instanton self-interaction, $V_{\text{inst}}=p_{0}-2\cos{\chi}$ in terms of the dimensionless real amplitude $\chi$ of the macroscopic wave function at $p_{0}=\pm 2$. Introducing the chemical potential $\mu$ and the internal pressure $p$ at zero temperature [see (35) and (43)], we classified stable, unstable and metastable states of matter using the $p-\mu$ phase diagram in the range $\chi\in[0;4.5\pi]$ in Fig. 3. The phase boundaries $\chi=a_{n}$ arise from the zeros of the spherical Bessel function, $j_{1}(a_{n})=0$ with $a_{1}=0$, defining intervals of stability $\chi\in(a_{2n};a_{2n+1})$ and instability $\chi\in(a_{2n-1};a_{2n})$ for $n\in\mathbb{N}$. Both the thermodynamic functions and the excitation spectrum confirm this partitioning, indicating that the squared sound speed $c^{2}_{s}(\chi)$ in (41) vanishes at $\chi=a_{n}$, is positive within stable bands, and negative within unstable ones. In particular, the small-amplitude regime at $\chi\in[0;\chi_{c}]$, where $\chi_{c}=a_{2}\simeq 4.493409$, is also shown to be unstable, which is consistent with earlier studies Nambo24 ; KT94 .
@@
We apply the BEC model (19)-(20), which takes into account self-gravity and quantum fluctuations, to nonperturbatively describe DM in dwarf galaxies over the range $\chi\in[0;2.5\pi]$, which spans exactly one stable and one unstable phase. After prescribing mathematically consistent initial conditions for the nonlinear spatial evolution, we fit a representative numerical solution to the rotation curve of the DM-dominated dwarf NGC 2366 Oh08 . The fit yields a dark boson mass $m\simeq 0.1171\times 10^{-22}~{}\text{eV}c^{-2}$, which is comparable to the mass in KMT . While NGC 2366 exhibits active star formation and ISM reionization, we employ its DM configuration solely as a prototypical case of our statistical ensemble of DM distributions in similar galactic settings.
Requiring that the stable phase forms the dense solitonic core of the DM halo, we compute thermodynamic functions for NGC 2366 and a family of hypothetical analogues with different core-to-halo mass ratios. By varying the partial pressure induced by quantum fluctuations $\Pi_{\rm q}$, we trace the resulting macroscopic characteristics in Fig. 10, which undergo a first-order phase transition for certain strengths of the gravitational parameter $A$, the only coupling constant in the model. The revealed jumplike behavior of the functions is attributed to the ambiguity in the quantum pressure $\Pi_{\rm q}$.
@@
Moreover, in the mass-radius diagram in Fig. 11(a) – which shows stable and unstable branches analogous to those of the axiton BZ19 ; KT94 ; ax5 – the DM configuration for NGC 2366 lies firmly on the stable branch, consistent with its existence in nature.
Thus, the thermodynamic signature provides a novel diagnostic for distinguishing stable and unstable DM configurations in galactic halos. Viewed ergodically, it provides an alternative to time-dependent approaches, and a direct comparison between the two deepens physical insight.

### v2

@@ III.2 Analyzing the macroscopic state of BEC DM
Having extracted the model parameters from the galaxy’s RC fit (e.g., NGC 2366), we then apply statistical analysis to probe the thermodynamic properties of BEC DM. Specifically, we focus on identifying the macroscopic state of DM in the galaxy. It depends on the fractions of the dense and rarefied phases of DM, and their change may be a phase transition. By mapping all configurations onto a diagram, the state of the galaxy would be represented by a single point there.
@@
When generating a set of solutions $\chi(\xi)$ for different values of $A$ and $\nu$ to support statistical analysis, we ensure that their shapes are similar to the reference solution shown in Fig. 4. To achieve this correspondence, varying $\nu$ at fixed $A$, we choose initial conditions from an “island” similar to that in Fig. 5 and impose certain conditions:
$\chi^{\prime}(\xi_{B})>0,\hskip 25.0pt\chi(\xi_{B})=\pi,$ (74)
which should hold simultaneously and define $\xi_{B}$.
Each wave function $\chi(\xi)$ in this set defines the chemical potential $\mu$, internal pressure $p$, and internal energy density $e$ according to
@@
Having designated the value of the wave function at the boundary as $\chi_{B}\equiv\chi(\xi_{B})$, and its derivative as $\chi^{\prime}_{B}\equiv\chi^{\prime}(\xi_{B})$, we proceed to the analysis of Eq. (79). Integrating by parts the trigonometric expressions, we arrive at
@@
Additionally, the mean pressure is evaluated in the units
$P_{0}\equiv\frac{\Gamma_{0}}{4\pi r_{0}^{3}}=\frac{\varepsilon_{0}}{v}=\frac{f_{\text{a}}^{2}}{2\hbar cr_{0}^{2}},$ (82)
yielding the numerical value $P_{0}=2.3383~{}\text{eV}\,\text{cm}^{-3}=3.7464\times 10^{-13}~{}\text{Pa}$ for the DM in NGC 2366.
@@
Denoting $\chi_{0}\equiv\chi(0)$ and expanding the set of previously introduced characteristics, we include the mean energy density ${\cal E}$ in Fig. 10(f):
@@
Thus, an increase in quantum fluctuations $\Pi_{\rm q}$ predictably leads to an expansion of the system, see Fig. 10(g). Consequently, the mean densities of particle number and internal energy naturally decrease, as shown in Fig. 10(d) and (f). Additionally, the enhanced fluctuations inhibit the increase of the central particle density, which is proportional to $\chi^{2}_{0}$, see Fig. 10(c). Moreover, the transition to the predominance of the rarefied phase is accompanied by an increase in chemical potential $\nu$, see Fig. 10(a). However, the decreasing near-linear part of the chemical potential $u$ in Fig. 10(b) indicates the unstable state.
@@
Valuable information can be also extracted from the mass-radius relation $M(R)$ by associating the total mass $M=4\pi r_{0}^{3}\rho_{0}N$ and the radius $R=r_{0}\xi_{B}$ with dimensionless quantities in Fig. 11(a). Although the number $N$ in the case of the unstable rarefied phase dominance tends to increase rapidly with radius $\xi_{B}>8$, we note the presence of a minimum $N_{\min}$ for relatively dense and compact objects at $\xi_{B}<8$. The existence of a minimum point separating the stable and unstable branches in our approach reproduces the predictions for the axitons BZ19 ; KT94 ; ax5 .
@@
The BEC DM state of NGC 2366, marked by a red dot in Figs. 10 and 11 at $\Pi_{\rm q}\simeq 5.852$, is characterized by the presence of $x\simeq 19.2\%$ of the dense DM in Fig. 10(h), which is concentrated in $4.7\%$ of the volume. According to our model, it belongs to the (blue) continuous trajectory, which is classified as supercritical fluid with respect to the variations in $(A,\Pi_{\rm q})$. However, this conclusion relies on spatially averaged values and overlooks the inhomogeneous DM distribution, as evidenced in Fig. 4 and 9. In addition, the phase trajectory of a macroscopic system can be significantly altered by random or external influences. In fact, BEC DM of NGC 2366 exhibits a structured, fruitlike composition, consisting of a dense core, sparse pulp, and crust. These zones are defined by distinct radial intervals: 1) $r\leq 2.83$ kpc, including a central region for $r\leq 0.94$ kpc with a density ratio $\chi_{0}^{2}/\sigma\simeq 6.04$; 2) $2.83\leq r\leq 5.55$ kpc; 3) $5.55\leq r\leq 7.79$ kpc, representing the outer layer in the model.
@@
On the other hand, a sharp decrease in $\Omega$ signals a fall into an unstable configuration.
@@
In Fig. 13(b), it is shown that the impact of surface gravity, expressed as ${\rm g}=AN/\xi^{2}_{B}$, also confirms the presence of a first-order phase transition in the model, although ${\rm g}$ is not a thermodynamic function (potential) used to extract critical parameters. We see that its value becomes larger as $x$ increases.
@@ IV Conclusion
We developed a statistical approach to the stationary nonrelativistic BEC DM model with instanton self-interaction, $V_{\text{inst}}=p_{0}-2\cos{\chi}$ in terms of the dimensionless real amplitude $\chi$ of the macroscopic wave function at $p_{0}=\pm 2$. Introducing the chemical potential $\mu$ and the internal pressure $p$ at zero temperature [see (35) and (43)], we classified stable, unstable, and metastable states of matter using the $p-\mu$ phase diagram in the range $\chi\in[0;4.5\pi]$ in Fig. 3. The phase boundaries $\chi=a_{n}$ arise from the zeros of the spherical Bessel function, $j_{1}(a_{n})=0$ with $a_{1}=0$, defining intervals of stability $\chi\in(a_{2n};a_{2n+1})$ and instability $\chi\in(a_{2n-1};a_{2n})$ for $n\in\mathbb{N}$. Both the thermodynamic functions and the excitation spectrum confirm this partitioning, indicating that the squared sound speed $c^{2}_{s}(\chi)$ in (41) vanishes at $\chi=a_{n}$, is positive within stable bands, and negative within unstable ones. In particular, the small-amplitude regime at $\chi\in[0;\chi_{c}]$, where $\chi_{c}=a_{2}\simeq 4.493409$, is also shown to be unstable, which is consistent with earlier studies Nambo24 ; KT94 .
@@
We apply the BEC model (19)-(20), which takes into account self-gravity and quantum fluctuations, to nonperturbatively describe DM in dwarf galaxies over the range $\chi\in[0;2.5\pi]$, which spans exactly one stable and one unstable phase. After prescribing mathematically consistent initial conditions for the nonlinear spatial evolution (66), we fit a representative $s$-wave numerical solution to the rotation curve of the DM-dominated dwarf NGC 2366 Oh08 ; Bl08 . The fit yields a dark boson mass $m\simeq 0.1171\times 10^{-22}~{}\text{eV}c^{-2}$, which is comparable to the mass in KMT . While NGC 2366 exhibits active star formation and ISM reionization Hun01 , we use its DM configuration solely as a prototypical case of our statistical ensemble of DM distributions in similar galactic settings.
By requiring the stable phase to form the dense solitonic core of the DM halo, we compute mean thermodynamic functions for NGC 2366 and a family of hypothetical analogs under similar boundary conditions (66) and (74), but with different core-to-halo mass ratios. By varying the partial pressure induced by quantum fluctuations $\Pi_{\rm q}$, we trace the resulting macroscopic characteristics in Fig. 10, which undergo a first-order phase transition for certain strengths of the gravitational parameter $A$, the only coupling constant in the model. The revealed jumplike behavior of the functions is attributed to the ambiguity in the quantum pressure $\Pi_{\rm q}$. To detect the phase transition, one can also use the compression $\Pi_{\nu}$ from GKN20 with similar properties and even the surface gravity value, see Fig. 13(b).
@@
Moreover, in the mass-radius diagram in Fig. 11(a) – which shows stable and unstable branches analogous to those of the axitons BZ19 ; KT94 ; ax5 – the DM configuration for NGC 2366 lies firmly on the stable branch, consistent with the observation. On the other hand, the internal mechanisms of the model are effective in addressing the problem of gravitational collapse Khl of the BEC DM halo, see e.g. Guzman ; H2019 ; H2014 .
Thus, the thermodynamic signature provides a novel diagnostic for distinguishing stable and unstable DM configurations in galactic halos. Viewed from an ergodic perspective, it offers an alternative to time-dependent approaches, and a direct comparison between the two deepens the understanding of the DM halo evolution.
Of course, the physical picture goes beyond macroscopic properties. In this regard, we mention at least the ability of DM particles in phase-1, i.e. at $\chi_{c}<\chi\lesssim 2\pi$, to form composites – two-particle dimers – during resonant scattering, which follows from GN23 .